# Localization phenomena in the random XXZ spin chain

### v1

@@ 2. Main result and technical steps
We fix $\Delta>1$ and $\lambda>0$. Let $\left\lvert S\right\rvert$ stand for the cardinality of $S\subset\mathbb{Z}$. Given $\Lambda\subset\mathbb{Z}$, we let $\mathcal{P}_{f}(\Lambda)=\left\{\mathbf{x}\subset\Lambda,\left\lvert\mathbf{x}\right\rvert<\infty\right\}$ be the collection of finite subsets of $\Lambda$, and let $\mathcal{P}_{N}(\Lambda)=\left\{\mathbf{x}\subset\Lambda,\left\lvert\mathbf{x}\right\rvert=N\right\}$ for $N\in\mathbb{N}^{0}$ be the subset of $\mathcal{P}_{f}(\Lambda)$ consisting of sets with cardinality $N$. We will use notation $\mathcal{P}_{+}(\Lambda)$ for the set $\mathcal{P}_{f}(\Lambda)\setminus\left\{\emptyset\right\}$.
Given $\Lambda\subset\mathbb{Z}$, let $H^{\Lambda}$ be the Hamiltonian given in (1.1), and consider the canonical orthonormal basis $\Phi_{\Lambda}=\left\{\phi_{\mathbf{x}}\right\}_{\mathbf{x}\in\mathcal{P}_{f}(\Lambda)}$ for $\mathcal{H}_{\Lambda}$, where
@@
Note that $\phi_{\emptyset}$, the vacuum state, is an eigenvector for $H^{\Lambda}$ with the simple eigenvalue 00. (It is the ground state for $H^{\Lambda}$, as we shall see later.) Note also that $\phi_{\mathbf{x}}^{\Lambda}=\phi_{\mathbf{x}}^{S}\otimes\phi_{\emptyset}^{\Lambda\setminus S}$ for $\mathbf{x}\subset S\subset\Lambda$. (We will suppress the $\Lambda$ dependence from $\phi_{\mathbf{x}}$ for ease of notation when it is clear from the context.) $\Phi_{\Lambda}$ can be decomposed as the disjoint union
$\Phi_{\Lambda}=\bigcup_{N=0}^{\left\lvert\Lambda\right\rvert}\Phi_{\Lambda}^{\left(N\right)},\mbox{ where }\Phi_{\Lambda}^{\left(N\right)}=\left\{\phi_{\mathbf{x}},\mathbf{x}\in\mathcal{P}_{N}(\Lambda)\right\}.$ (2.2)
If $\mathbf{x}\in\mathcal{P}_{N}(\Lambda)$ with $N\geq 1$, we identify $\mathbf{x}$ with $(x_{1},\ldots,x_{N})\in\Lambda^{N}$, where $x_{1}<x_{2}<\ldots<x_{N}$, and set $\left\lvert x\right\rvert_{1}=\sum_{i=1}^{N}\left\lvert x_{i}\right\rvert$ and $\left\lvert\mathbf{x}\right\rvert_{2}=\left(\sum_{i=1}^{N}x_{i}^{2}\right)^{\frac{1}{2}}$.
@@
Note $\mathcal{N}_{i}=P_{-}^{\left\{i\right\}}$ for all $i\in\mathbb{Z}$.
For a unit vector $\varphi\in\mathcal{H}_{\Lambda}$, we denote by $\pi_{\varphi}$ the orthogonal projection onto $\varphi$. If $\mathbf{u}\in\mathcal{P}_{f}(\Lambda)$, we write $\pi_{\mathbf{u}}=\pi_{{\phi_{\mathbf{u}}}}$. We have
@@
More generally, for an operator $T$ on $\mathcal{H}_{\Lambda}$ we denote by $T({\bf x},{\bf y})$ its matrix elements with respect to the standard basis, i.e.,
@@
We equip $\mathbb{Z}$ with the usual graph distance $d_{\mathbb{Z}}(x,y)=\left\lvert x-y\right\rvert$ for $x,y\in\mathbb{Z}$. We will consider $\Lambda\subset\mathbb{Z}$ as a subgraph of $\mathbb{Z}$, and denote by $\operatorname{dist}_{\Lambda}(\cdot,\cdot)$ the graph distance in $\Lambda$, which can be infinite if $\Lambda$ is not a connected subset of $\mathbb{Z}$. Given $S\subset\Lambda\subset\mathbb{Z}$ and $p\in\mathbb{N}^{0}=\left\{0\right\}\cup\mathbb{N}$, we set
${[S]^{\Lambda}_{p}}$ =$=\left\{x\in\Lambda:\operatorname{dist}_{\Lambda}\left(x,S\right)\leq p\right\},$ (2.8)
$\partial_{ex}^{\Lambda}S$ =$=\left\{x\in\Lambda:\ d_{\Lambda}\left(x,S\right)=1\right\}=[S^{\Lambda}]_{1}\setminus S,$
$\partial_{in}^{\Lambda}S$ =$=\left\{x\in\Lambda:\ \operatorname{dist}_{\Lambda}\left(x,\Lambda\setminus S\right)=1\right\},$
@@
$\widetilde{d}_{H}^{\Lambda}(\mathbf{x},\mathbf{y})=\begin{cases}d_{H}^{\Lambda}(\mathbf{x},\mathbf{y})&\quad\text{if}\quad\left\lvert\mathbf{x}\right\rvert=\left\lvert\mathbf{y}\right\rvert\\ \infty&\quad\text{otherwise}\end{cases}.$ (2.12)
@@
We will denote by $\mathbb{E}_{\Lambda}$ the expectation with respect to the random variables $\left\{\omega_{i}\right\}_{i\in\Lambda}$. In this paper we will use generic constants $C,c$, etc., whose values will be allowed to change from line to line, even in the same line in a displayed equation. These constants will, in general, depend on the fixed parameters of the model such as $\mu$, $\Delta$, $\lambda$, and on the fractional moment exponent $s$, but (critically) they will be volume-independent. We will not indicate the dependence on the fixed parameters and on $s$, but, when necessary, we will indicate the dependence of a constant on other parameters, say $q,N,\dots$, explicitly by writing the constant as $C_{q},C_{q,N},\ldots$. If we write $C_{q}$, it does not depend on $N$. These constants can always be estimated from the arguments, but we will not track the changes to avoid complicating the arguments. We will use $C$ to indicate that the constant should be sufficiently large for a bound to hold, and $c$ to indicate that the constant should be sufficiently small, but still requiring $c>0$ We generally use the same $C$ and $c$ for different constants in the same equation.
@@
The result above is not vacuous as the spectrum $\sigma(H^{\mathbb{Z}})=\left\{0\right\}\cup[1-\frac{1}{\Delta},\infty)$ with probability one. (See, e.g., the discussion in [EKS1]). Note also that 00 is a simple eigenvalue.
The key input for proving Theorem 2.1 is an immediate corollary to [EK22, Theorem 2.4], which we now state.
@@
Fix parameters $\Delta_{0}>1$ and $\lambda_{0}>0$. Let $q\in\frac{1}{2}\mathbb{N}^{0}$ and $s\in(0,\frac{1}{3})$. Then there exists a constant $Y$ (which depends on $\Delta_{0}\,,\lambda_{0}$, $\mu$, $q$, and $s$) such that, for all $\Delta\geq\Delta_{0}$ and $\lambda\geq\lambda_{0}$ satisfying $\lambda\Delta^{2}\geq Y$ the following holds: For all finite $D\subset\mathbb{Z}$ we have
@@
We proved a slightly stronger result in [EK22, Theorem 2.4], where it is shown that under the hypotheses of the theorem there exists a constant $Y$ (which depends on $\Delta_{0}$, $\lambda_{0}$, $\mu$, $s$, and $q$) such that, for all $\Delta\geq\Delta_{0}$ and $\lambda\geq\lambda_{0}$ satisfying $\lambda\Delta^{2}\geq Y$, for all $D\subset\mathbb{Z}$ finite we have
@@
for all $A\subset B\subset D$ with $A$ connected in $D$. ([EK22, Theorem 2.4] is stated and proved for real energies in the intervals $(-\infty,k+\tfrac{3}{4}]$, where $k\in\mathbb{N}^{0}$. The proof is also valid for complex energies $z$ with $\operatorname{Re}z\leq(k+\tfrac{3}{4})\left(1-\frac{1}{\Delta}\right)$, with the same constants. The above result follows.)
@@
and hence (2.17) follows from (2.18) as $d_{D}\left(\left\{x\right\},\Lambda\setminus{[x]^{D}_{r-1}}\right)\geq r$. ∎
We now state our main technical result, Theorem 2.4 bellow. But first we need to introduce some additional notation and observations.
Let $S\subset\mathbb{Z}$. Given an energy interval $I$, we set $\sigma_{I}(H^{S})=\sigma(H^{S})\cap I$. If $\nu\in\mathbb{R}$, we set $\pi^{S}_{\nu}=\chi_{\left\{\nu\right\}}(H^{S})$, the spectral projection of $H^{S}$ on the set $\left\{\nu\right\}$.
The eigencorrelator $\mathcal{Q}^{S}_{I}$ for $H^{S}$ in the energy interval $I$ is given by
@@
We only consider ${\bf x},{\bf y}\in\mathcal{P}_{+}(\Lambda)$ because $G_{z}^{\Lambda}(\emptyset,\emptyset)=-\tfrac{1}{z}$ for $z\neq 0$ and and $G_{z}^{\Lambda}(\emptyset,\mathbf{x})=0$ for ${\bf x}\in\mathcal{P}_{+}(\Lambda)$. More generally, given a bounded Borel measurable function $f$, we have $f(H^{\Lambda})(\emptyset,\emptyset)=f(0)$ and $f(H^{\Lambda})(\emptyset,\mathbf{x})=0$ for ${\bf x}\in\mathcal{P}_{+}(\Lambda)$.
@@
The input in the theorem, the estimate (2.22) (the finite volumes criterion), allows for volume dependence, whereas the output (2.23) completely suppresses this dependence. From the technical point of view, this is one of the delicate points in the analysis, and the supression of the volume dependence is a crucial step in proving Theorem 2.4. In addition, the output of the theorem is also valid for infinite subsets $\Lambda$ of $\mathbb{Z}$.
The proof of Theorem 2.4, given in Section 4, proceeds by induction over $q\in\frac{1}{2}\mathbb{N}^{0}$, with constants $C_{q}$ and $c_{q}$ in (2.23) that deteriorate with $q$, rending the method unpractical beyond fixed energy intervals. Similarly to the situation with random Schrödinger operators in dimension higher than one, it is not clear whether this restriction is a technical shortcoming or a feature (i.e., there is a phase transition for high energies for this model). There is no consensus among physicists whether such phase transition occurs or not in the infinite volume systems.
Theorem 2.1 follows immediately from Theorems  2.3, 2.4, and Theorem 2.6 bellow.
@@
Moreover, there exists an event $\mathcal{E}$, with $\mathbb{P}_{\mathbb{Z}}(\mathcal{E})=1$, such that for $\omega\in\mathcal{E}$ the spectrum of $H^{\mathbb{Z}}$ in $I_{\leq q}$ is pure point, and if $\psi_{\omega}$ is an eigenfunction of $H^{\mathbb{Z}}$ with corresponding eigenvalue in $I_{q}$, so $\psi\in\mathcal{H}_{\mathbb{Z}}^{N_{\psi}}$ for some $N_{\psi}\in\mathbb{N}$, it decays exponentially in the following sense:
@@
Theorem 2.6 is proven in Section 5, where it is derived from [AW2, Theorem 4.1].
The rest of this paper is organized as follows: In Section 3 we introduce notation and collect some basic properties of the XXZ spin chain that are used in our arguments. We prove Theorems 2.4 and 2.6 in Sections 4 and 5, respectively. In Appendix A we provide a more detailed discussion of localization types for single-particle and many-body systems. Appendix B provides bounds on exponential sums that will be encountered throughout the paper. In Appendix C we discuss useful properties of the so-called filter function that appears in the proof of Theorem 2.4.
@@ 3. Basic features of the XXZ spin chain
When working with a fixed $\Lambda\subset\mathbb{Z}$, we write $K^{c}=\Lambda\setminus K$ for $K\subset\Lambda$.
We note that $\mathcal{N}_{i}$ is the projection onto the spin-down state (also called the local number operator) at site $i$. Given $S\subset\Lambda$, $\mathcal{N}^{S}=\sum_{i\in S}\mathcal{N}_{i}$ is the total (spin-down) number operator in $S$. The total number operator $\mathcal{N}^{\Lambda}$ has eigenvalues $0,1,2,\ldots,\left\lvert\Lambda\right\rvert$. We set $\mathcal{H}_{\Lambda}^{\left(N\right)}=\operatorname{Ran}{\chi_{\left\{N\right\}}(\mathcal{N}^{\Lambda})}$, obtaining the Hilbert space decomposition $\mathcal{H}_{\Lambda}=\bigoplus_{N=0}^{\left\lvert\Lambda\right\rvert}\mathcal{H}_{\Lambda}^{\left(N\right)}$.
@@
We note that the operators $\left\{\mathcal{N}_{i}\right\}$ (and thus $\mathcal{W}^{\Lambda}$ and $V_{\omega}$) are diagonal in the canonical basis: $\mathcal{N}_{i}\phi_{\bf x}=\phi_{\bf x}$ if $i\in{\bf x}$ and 00 otherwise. $\mathcal{W}^{\Lambda}$ is the number of clusters operator: $\mathcal{W}^{\Lambda}\phi_{\mathbf{x}}=W^{\Lambda}_{\mathbf{x}}\phi_{\mathbf{x}}$ for $\mathbf{x}\subset\Lambda$ finite, where $W^{\Lambda}_{\mathbf{x}}$ is the number of clusters (connected components) of $\mathbf{x}$ as a subset of $\Lambda$, so $\sigma\left(\mathcal{W}^{\Lambda}\right)\subset\left\{0,1,2,\ldots,\left\lvert\Lambda\right\rvert\right\}$. $V^{\Lambda}_{\omega}$ is the random potential:
$V^{\Lambda}_{\omega}\phi_{\mathbf{x}}=V_{\omega}(\mathbf{x})\phi_{\mathbf{x}}\;\ \text{for}\;\ \mathbf{x}\subset\Lambda\;\text{finite},\;\text{where}\;V_{\omega}(\mathbf{x})=\left(\sum_{i\in\mathbf{x}}\omega_{i}\right).$ (3.2)
An important feature of the XXZ Hamiltonian $H^{\Lambda}$ is total particle number (or magnetization) preservation: the operators $\boldsymbol{\Delta}^{\Lambda}$and the total number of particles operator $\mathcal{N}^{\Lambda}=\sum_{i\in\Lambda}\mathcal{N}_{i}$ commute (all bounded functions of these operators commute), and hence $H^{\Lambda}$ and $\mathcal{N}^{\Lambda}$ also commute. If $\Lambda$ is finite, this is equivalent to
$=-\tfrac{1}{2\Delta}[\boldsymbol{\Delta}^{\Lambda},\mathcal{N}^{\Lambda}]=0.$ (3.3)
It can be verified (e.g., [EK22]) that

### v3

@@ 2. Main result and technical steps
We let $\left\lvert S\right\rvert$ stand for the cardinality of $S\subset\mathbb{Z}$. Given $\Lambda\subset\mathbb{Z}$, we let
$\mathcal{P}_{f}(\Lambda)=\left\{\mathbf{x}\subset\Lambda:\left\lvert\mathbf{x}\right\rvert$$\infty\right\}$
be the collection of finite subsets of $\Lambda$, and let
$\mathcal{P}_{N}(\Lambda)=\left\{\mathbf{x}\subset\Lambda:\left\lvert\mathbf{x}\right\rvert=N\right\}$
for $N\in\mathbb{N}^{0}=\mathbb{N}\cup\left\{0\right\}$ be the subset of $\mathcal{P}_{f}(\Lambda)$ consisting of sets with cardinality $N$. We use the notation $\mathcal{P}_{+}(\Lambda)=\mathcal{P}_{f}(\Lambda)\setminus\left\{\emptyset\right\}$. We also set $\left\lvert\mathbf{x}\right\rvert_{1}=\sum_{x\in\mathbf{x}}\left\lvert x\right\rvert$ and $\left\lvert\mathbf{x}\right\rvert_{2}=\left(\sum_{x\in\mathbf{x}}\left\lvert x\right\rvert^{2}\right)^{\frac{1}{2}}$ for $\mathbf{x}\in\mathcal{P}_{f}(\mathbb{Z})$.
We fix $\Delta>1$ and $\lambda>0$. Given $\Lambda\subset\mathbb{Z}$, let $H^{\Lambda}$ be the Hamiltonian given in (1.1), and consider the canonical orthonormal basis $\Phi_{\Lambda}=\left\{\phi_{\mathbf{x}}\right\}_{\mathbf{x}\in\mathcal{P}_{f}(\Lambda)}$ for $\mathcal{H}_{\Lambda}$, where
@@
Note that $\phi_{\emptyset}$, the vacuum state, is an eigenvector for $H^{\Lambda}$ with the simple eigenvalue 0. (It is the ground state for $H^{\Lambda}$, as we shall see later.) Note also that $\phi_{\mathbf{x}}^{\Lambda}=\phi_{\mathbf{x}}^{S}\otimes\phi_{\emptyset}^{\Lambda\setminus S}$ for $\mathbf{x}\subset S\subset\Lambda$. (We will suppress the $\Lambda$ dependence from $\phi_{\mathbf{x}}$ for ease of notation when it is clear from the context.) $\Phi_{\Lambda}$ can be decomposed into the disjoint union
$\Phi_{\Lambda}=\bigcup_{N=0}^{\left\lvert\Lambda\right\rvert}\Phi_{\Lambda}^{\left(N\right)},\mbox{ where }\Phi_{\Lambda}^{\left(N\right)}=\left\{\phi_{\mathbf{x}}:\mathbf{x}\in\mathcal{P}_{N}(\Lambda)\right\}.$ (2.2)
@@
Note that $\mathcal{N}_{i}=P_{-}^{\left\{i\right\}}$ for all $i\in\mathbb{Z}$.
For a unit vector $\varphi\in\mathcal{H}_{\Lambda}$, we denote by $\pi_{\varphi}$ the orthogonal projection onto $\mathbb{C}\varphi$. If $\mathbf{u}\in\mathcal{P}_{f}(\Lambda)$, we write $\pi_{\mathbf{u}}=\pi_{{\phi_{\mathbf{u}}}}$. We have
@@
(For real valued $z$, $G^{\Lambda}_{z}({\bf x},{\bf y})$ is understood here and below as $G^{\Lambda}_{z+i0}({\bf x},{\bf y})$.) More generally, for an operator $T$ on $\mathcal{H}_{\Lambda}$ we denote by $T({\bf x},{\bf y})$ its matrix elements with respect to the standard basis, i.e.,
@@
We equip $\mathbb{Z}$ with the usual graph distance $d_{\mathbb{Z}}(x,y)=\left\lvert x-y\right\rvert$ for $x,y\in\mathbb{Z}$. We will consider $\Lambda\subset\mathbb{Z}$ as a subgraph of $\mathbb{Z}$, and denote by $\operatorname{dist}_{\Lambda}(\cdot,\cdot)$ the graph distance in $\Lambda$, which can be infinite if $\Lambda$ is not a connected subset of $\mathbb{Z}$. Given $S\subset\Lambda\subset\mathbb{Z}$ and $p\in\mathbb{N}^{0}$, we set
${[S]^{\Lambda}_{p}}$ $=\left\{x\in\Lambda:\operatorname{dist}_{\Lambda}\left(x,S\right)\leq p\right\},$ (2.8)
$\partial_{ex}^{\Lambda}S$ $=\left\{x\in\Lambda:\ d_{\Lambda}\left(x,S\right)=1\right\}=[S^{\Lambda}]_{1}\setminus S,$
$\partial_{in}^{\Lambda}S$ $=\left\{x\in\Lambda:\ \operatorname{dist}_{\Lambda}\left(x,\Lambda\setminus S\right)=1\right\},$
@@
[…]
@@
We will denote by $\mathbb{E}_{\Lambda}$ the expectation with respect to the random variables $\left\{\omega_{i}\right\}_{i\in\Lambda}$. In this paper we will use generic constants $C,c$, etc., whose values will be allowed to change from line to line, even in a multi-line displayed equation. These constants will, in general, depend on the fixed parameters of the model such as $\mu$, $\Delta$, $\lambda$, and on the fractional moment exponent $s$, but (critically) they will be volume-independent. We will not indicate the dependence on the fixed parameters and $s$, but, when necessary, we will indicate the dependence of a constant on other parameters, say $q,N,\dots$, explicitly by writing the constant as $C_{q},C_{q,N},\ldots$. If we write $C_{q}$, it does not depend on $N$. These constants can always be estimated from the arguments, but we will not track the changes to avoid complicating the arguments. We will use $C$ to indicate that the constant should be sufficiently large for a bound to hold, and $c$ to indicate that the constant should be sufficiently small, but still requiring $c>0$ . We generally use the same $C$ and $c$ for different constants in the same equation.
@@
The result above is not as vacuous as the spectrum $\sigma(H^{\mathbb{Z}})=\left\{0\right\}\cup[1-\frac{1}{\Delta},\infty)$ with probability one. (See, e.g., the discussion in [15]). Note also that 0 is a simple eigenvalue.
The key input for proving Theorem 2.1 is an immediate corollary to [12, Theorem 2.4], which we now state.
@@
Fix the parameters $\Delta_{0}>1$ and $\lambda_{0}>0$. Let $q\in\frac{1}{2}\mathbb{N}^{0}$ and $s\in(0,\frac{1}{3})$. Then there exists a constant $Y$ (which depends on $\Delta_{0}\,,\lambda_{0}$, $\mu$, $q$, and $s$) such that, for all $\Delta\geq\Delta_{0}$ and $\lambda\geq\lambda_{0}$ satisfying $\lambda\Delta^{2}\geq Y$ the following holds: For all finite $D\subset\mathbb{Z}$ we have
@@
We proved a stronger result in [12, Theorem 2.4], where it is shown that under the hypotheses of the theorem there exists a constant $Y$ (which depends on $\Delta_{0}$, $\lambda_{0}$, $\mu$, $s$, and $q$) such that, for all $\Delta\geq\Delta_{0}$ and $\lambda\geq\lambda_{0}$ satisfying $\lambda\Delta^{2}\geq Y$, for all $D\subset\mathbb{Z}$ finite we have
@@
for all $A\subset B\subset D$ with $A$ connected in $D$. ([12, Theorem 2.4] is stated and proved for real energies in the intervals $(-\infty,k+\tfrac{3}{4}]$, where $k\in\mathbb{N}^{0}$. The proof is also valid for complex energies $z$ with $\operatorname{Re}z\leq(k+\tfrac{3}{4})\left(1-\frac{1}{\Delta}\right)$, with the same constants. The above result follows.)
@@
hence (2.17) follows from (2.18) as $d_{D}\left(\left\{x\right\},\Lambda\setminus{[x]^{D}_{r-1}}\right)\geq r$. ∎
We now state our main technical result, Theorem 2.4 below. But first we need to introduce some additional notation and observations.
Let $S\subset\mathbb{Z}$. Given an energy interval $I$, we set $\sigma_{I}(H^{S})=\sigma(H^{S})\cap I$. If $\nu\in\mathbb{R}$, we set $\pi^{S}_{\nu}=\chi_{\left\{\nu\right\}}(H^{S})$, the spectral projection of $H^{S}$ on the set $\left\{\nu\right\}$. The eigencorrelator $\mathcal{Q}^{S}_{I}$ for $H^{S}$ in the energy interval $I$ is given by
@@
We only consider ${\bf x},{\bf y}\in\mathcal{P}_{+}(\Lambda)$ because $G_{z}^{\Lambda}(\emptyset,\emptyset)=-\tfrac{1}{z}$ for $z\neq 0$ and $G_{z}^{\Lambda}(\emptyset,\mathbf{x})=0$ for ${\bf x}\in\mathcal{P}_{+}(\Lambda)$. More generally, given a bounded Borel measurable function $f$, we have $f(H^{\Lambda})(\emptyset,\emptyset)=f(0)$ and $f(H^{\Lambda})(\emptyset,\mathbf{x})=0$ for ${\bf x}\in\mathcal{P}_{+}(\Lambda)$.
@@
The input in the theorem, the estimate (2.22) (the finite volumes criterion), allows for volume dependence, whereas the output (2.23) completely suppresses this dependence. From a technical point of view, this is one of the delicate points in the analysis, and the suppression of volume dependence is a crucial step in proving Theorem 2.4. Note that the output (2.23) is also valid for infinite subsets $\Lambda$ of $\mathbb{Z}$.
The proof of Theorem 2.4, presented in Section 4, proceeds by induction over $q\in\frac{1}{2}\mathbb{N}^{0}$, with constants $C_{q}$ and $c_{q}$ in (2.23) that deteriorate with $q$, rendering the method unpractical beyond fixed energy intervals. Similarly to the case of random Schrödinger operators in dimension higher than one, it is not clear whether this restriction is a technical shortcoming or a feature (i.e., there is a phase transition for high energies for this model). There is no consensus among physicists on whether such phase transition occurs or not in the infinite volume system.
Theorem 2.1 follows immediately from Theorem  2.3, Theorem 2.4, and Theorem 2.6 below.
@@
Moreover, there exists an event $\mathcal{E}$, with $\mathbb{P}_{\mathbb{Z}}(\mathcal{E})=1$, such that for $\omega\in\mathcal{E}$ the spectrum of $H^{\mathbb{Z}}$ in $I_{\leq q}$ is pure point, and if $\psi_{\omega}$ is an eigenfunction of $H^{\mathbb{Z}}$ with the corresponding eigenvalue in $I_{q}$, so $\psi\in\mathcal{H}_{\mathbb{Z}}^{N_{\psi}}$ for some $N_{\psi}\in\mathbb{N}$, it decays exponentially in the following sense:
@@
Theorem 2.6 is proven in Section 5, where it is derived from [3, Theorem 4.1].
The rest of this paper is organized as follows: In Section 3 we introduce notation and collect some basic properties of the XXZ spin chain. We prove Theorems 2.4 and 2.6 in Sections 4 and 5, respectively. In Appendix A we provide a more detailed discussion of localization types for single-particle and many-body systems. Appendix B provides bounds for the exponential sums encountered throughout the paper. In Appendix C we discuss key properties of the filter function that appears in the proof of Theorem 2.4.
@@ 3. Basic features of the XXZ spin chain
Let $\Lambda\subset\mathbb{Z}$. When working with a fixed $\Lambda$, we write $K^{c}=\Lambda\setminus K$ for $K\subset\Lambda$. If $\mathbf{x}\subset\Lambda$ and $S\subset\Lambda$, we write $\mathbf{x}_{S}=\mathbf{x}\cap S$. If $P$ is an orthogonal projection, we write $\bar{P}=\mathds{1}-P$.
Recall that $\mathcal{N}_{i}$ is the orthogonal projection onto the spin-down state (called the local number operator) at the site $i$. Given $S\subset\Lambda$, $\mathcal{N}^{S}=\sum_{i\in S}\mathcal{N}_{i}$ is the total (spin-down) number operator in $S$. The total number operator $\mathcal{N}^{\Lambda}$ has eigenvalues $0,1,2,\ldots,\left\lvert\Lambda\right\rvert$. We set $\mathcal{H}_{\Lambda}^{\left(N\right)}=\operatorname{Ran}{\chi_{\left\{N\right\}}(\mathcal{N}^{\Lambda})}$, obtaining the Hilbert space decomposition $\mathcal{H}_{\Lambda}=\bigoplus_{N=0}^{\left\lvert\Lambda\right\rvert}\mathcal{H}_{\Lambda}^{\left(N\right)}$.
@@
Note that the operators $\left\{\mathcal{N}_{i}\right\}$ (and thus $\mathcal{W}^{\Lambda}$ and $V_{\omega}$) are diagonal in the canonical basis: $\mathcal{N}_{i}\phi_{\bf x}=\phi_{\bf x}$ if $i\in{\bf x}$ and 0 otherwise. $\mathcal{W}^{\Lambda}$ is the number of clusters operator: $\mathcal{W}^{\Lambda}\phi_{\mathbf{x}}=W^{\Lambda}_{\mathbf{x}}\phi_{\mathbf{x}}$ for $\mathbf{x}\subset\Lambda$ finite, where $W^{\Lambda}_{\mathbf{x}}$ is the number of clusters (connected components) of $\mathbf{x}$ as a subset of $\Lambda$, so $\sigma\left(\mathcal{W}^{\Lambda}\right)\subset\left\{0,1,2,\ldots,\left\lvert\Lambda\right\rvert\right\}$. $V^{\Lambda}_{\omega}$ is the random potential:
$V^{\Lambda}_{\omega}\phi_{\mathbf{x}}=V_{\omega}(\mathbf{x})\phi_{\mathbf{x}}\;\ \text{for}\;\ \mathbf{x}\in\mathcal{P}_{f}(\Lambda),\quad\text{where}\quad V_{\omega}(\mathbf{x})=\left(\sum_{i\in\mathbf{x}}\omega_{i}\right).$ (3.2)
An important feature of the XXZ Hamiltonian $H^{\Lambda}$ is the conservation of the total particle number (or magnetization): the operators $\boldsymbol{\Delta}^{\Lambda}$ and the total number of particles operator $\mathcal{N}^{\Lambda}=\sum_{i\in\Lambda}\mathcal{N}_{i}$ commute (i.e., all bounded functions of these operators commute), and hence $H^{\Lambda}$ and $\mathcal{N}^{\Lambda}$ also commute. If $\Lambda$ is finite, this is equivalent to
$\ [H^{\Lambda},\mathcal{N}^{\Lambda}]=-\tfrac{1}{2\Delta}[\boldsymbol{\Delta}^{\Lambda},\mathcal{N}^{\Lambda}]=0.$ (3.3)
It can be verified (e.g., [12]) that